OpenStudy (anonymous):

What is the quotient of the rational expression below? x^2 - 25 / x- 11 / x^2 + 10x + 25/ 4x -44 Help Please!!!!! Correct answer gets a MEDAL!!!!!!!!

OpenStudy (anonymous):

Just in case you didn't understand

OpenStudy (triciaal):

to divide by a fraction invert and multiply. to simplify find factors that can cancel out x^2-25 = difference of 2 squares = (x + 5)(x-5) 4x-44 = 4(x - 11) x^2 + 10 x + 25 = (x+5)(x+5) rewrite the original using the factored form then simplify

OpenStudy (anonymous):

So the answer is A ? @triciaal

OpenStudy (triciaal):

no I have D

OpenStudy (triciaal):